# Recidivism and Peer Influence with LLM Text Embeddings in Low Security Correctional Facilities

### v1

@@ 5.2 Endogenous Network Formation
(c) Standard deviation of row means (Male)
(d) Standard deviation of row means (Female)
(e) Transitivity of clustering coefficient (Male)
(f) Transitivity of clustering coefficient (Female)
@@
We compare five estimators for four network models in terms of their ability to replicate observed salient features of the TC networks. These estimators/models are (1) tetra logit estimator (tet.logit) and the (2) joint fixed effect maximum likelihood estimator (jfe) in Graham (2017) that allow for single latent variable (additive) and covariates, 3) Spectral embedding in RDPG model (Athreya et al., 2017) which allows for multivariate latent factors but no covariates (rdpg), 4) MLE in latent space model that allows for both multiplicative and additive latent factors but no observed covariates Hoff et al. (2002) (lsm) and finally 5) MLE in latent space model from Ma et al. (2020) that allows for both multiplicative and additive latent factors and observed covariates (lsm-cov).
Among the above-listed estimators that allow for observed covariates for network formation, we use the time difference (in days) between the residents’ entry dates as a node-pair level covariate. The entry dates of residents can be assumed to be random and not explained by their observed or unobserved characteristics, since a complex function of court decisions and availability of beds in TCs determines them. However, the proximity of entry dates is an important determinant of whether residents will exchange messages since overlap in their time in the TC is a necessary condition to exchange messages. We do not use the proximity in exit dates or time in TC as a predictor of network formation, as the unobserved heterogeneity of residents likely impacts these variables, and they are not pre-determined to network formation.
We compare the estimators in terms of their ability to replicate three widely observed structural network properties: modularity (Girvan and Newman, 2002), standard deviation of row means, and transitivity or the clustering coefficient (Newman, 2018). The steps deployed for comparison are as follows. Let us consider the network data for the male unit. We implement each estimator on the observed male unit network to estimate the respective fitted models. Next, we simulate 200 networks using these fitted models (200 for each estimator listed above). The structural properties over these simulated networks are then compared with the true network to assess how well the simulated data can replicate the truth. Figure 8 displays the performance. We observe that the latent space model that allows for both multiplicative and additive latent factors, along with covariates (proximity in entry dates), outperforms all other models for every property in the male unit. For the female unit, it performs the best for the standard deviation of row means and transitivity of the clustering coefficient, and at least as good as the other estimators for modularity.
@@ 5.3 Peer Effects in Text
Personal Growth ALR Community Support ALR Rule Violations ALR
@@
This section provides our results on the peer effects model, where the outcome variables are the class probabilities generated by Zero-Shot classification. Our preferred specification involves estimating the network formation model using a latent space model, which allows for both multiplicative and additive latent variables and observed covariates (lsm-cov). This selection is based on the results obtained in the previous section, which show that the latent space model with covariates performs the best in replicating key structural properties of the true male and female unit networks.
@@
Personal Growth ALR Community Support ALR Rule Violations ALR
@@
Tables 3 and 4 provide the results for IV 2SLS in text class probabilities after adjusting for the unobserved heterogeneity in the outcome model (Algorithm 1) for females and males, respectively. The dimension of the multiplicative factor is chosen using data-driven cross-validation methods for networks in Li et al. (2020). The choice of data-driven dimensions for the multiplicative factors is 14 for the female and 19 for the male units, respectively. In addition, the model for network formation contains one additive nodal latent variable and an observed edge-level covariate that captures the difference in the residents’ entry dates.
First, we discuss the diagonal elements in Table 3, which inform us about the spillovers of peer probabilities on female residents for the same class. There is evidence of a peer effect in all three classes relative to disruptive conduct at the 5% significance level. Moreover, there is also evidence of peer spillovers from personal growth and community support categories on rule violations category. We can interpret these results as follows, for example. An individual is more likely to send messages of peer community support if their peers are also sending messages of community support. Table B2 presents the results when unobserved heterogeneity in the outcome model and network formation are not accounted for, for comparison.
@@
Overall, we find evidence of peer influence in the ALR of text class probabilities both when aggregated by the sender profiles and when aggregated by the receiver profiles. However, the magnitude and standard errors change considerably once we account for unobserved heterogeneity in outcome and network formation. Direct peer influence spillovers have a larger impact than indirect spillovers for both male and female units.
6 Conclusion
This paper deploys large language models to construct text embeddings from a large corpus of written exchanges between residents in low-security correctional facilities called Therapeutic Communities. These exchanges have substantial explanatory power in predicting recidivism up to three years after exiting the facility. The out-of-sample predictive accuracy is 30% higher with these text embeddings relative to using only pre-entry covariates in the prediction exercise. The paper further uses Zero-Shot classification to interpret these embeddings and learn about peer influences in sender and receiver profiles of these messages. To study peer effects on correlated multivariate LLM representations of language, we provide new methodological and theoretical results on an instrumental variable approach that allows for multiple dimensions of latent homophily, sparse network settings, and multivariate outcomes.
Acknowledgement and IRB
The Therapeutic Community data used for this paper were exempted from IRB review based on IRB Protocol 2007B0162 by The Ohio State University. SP’s research was partially funded by a grant from NSF DMS (2529302). We thank the Ohio Supercomputer Center (OSC) for providing us with computing resources through an academic account.
Appendix A Mathematical Appendix
@@ A.1 Proof of Lemma 1
We can easily see that the conditional expectation of the second part, conditioning on $\bm{U}_{i}$ is (adapting the arguments from Johnsson and Moon [2021])
@@
A.2 Proof of Theorem 1
We derive the asymptotic distribution of the 2SLS estimator $\hat{\boldsymbol{\beta}}_{v,\text{IV}}$ in three steps similar to the structure in Johnsson and Moon [2021].
@@ A.2 Proof of Theorem 1
Now, let ${\cal F}_{Ni}:=\sigma\left(\bm{X}_{N},\bm{A},\bm{u}_{s},\bm{e}_{s};\,s\leq i\right)$ be a filtration. Then, $T_{Ni}:=\frac{1}{\sqrt{N}}\tilde{\bm{k}}_{i}^{\prime}\bm{Q}\tilde{\bm{e}}_{i}$ is a martingale difference adapted to ${\cal F}_{Ni}$. For simplicity, let $\bm{\iota}_{i}^{E}:=\bm{Q}\tilde{\bm{e}}_{i}$, which are i.i.d from $(0_{3p},\bm{Q}\Sigma_{\tilde{E}}\bm{Q}^{\prime})$. We check the sufficient conditions for Corollary 3.1 of Hall and Heyde [2014] to apply the Martingale central limit theorem.
@@
    By Chen et al. [2020, equation 45 in Lemma A.8],
@@
A.3 Proof of additional lemmas
A.3.1 Proof of Lemma 3
This part of the proof follows similar logic as in Johnsson and Moon [2021]; however, since our terms are matrix-valued instead of vector-valued, our bounds involve the matrix spectral and Frobenius norms, as well as the matrix trace. We first prove the four results claimed in the lemma.
@@ A.3.1 Proof of Lemma 3
    under Assumption 3 using the results of Lemmas 2 and 7. Next,
@@ A.3.2 Proof of Lemma 4
A.4 Additional Technical Lemmas
@@
Let Assumption 3 hold. Under the conditions on RDPG model in Rubin-Delanchy et al. [2022] and sparsity parameter $\rho_{N}=\omega(\frac{\log^{4}N}{N})$, we have
@@
and under the conditions on the sparse LSM model with sparsity parameter $\omega_{N}=\omega(N^{-1/2})$ in Li et al. [2023], we have
@@
By the results on Adjacency Spectral Embedding in RDPG model in Rubin-Delanchy et al. [2022], the last term is less than
@@
Temporarily write $\bm{W}^{(i)}:=\bm{\phi}^{L_{N}}(\bm{u}_{i})\bm{\phi}^{L_{N}}(\bm{u}_{i})^{\prime}$. By Lemma 5 of Johnsson and Moon [2021], we have
@@
Appendix B Additional Tables and Figures
Table B1: Summary Statistics
@@ B. Male Unit
Table B2: Peer Effects in Text Class Probabilities (Female) IV 2SLS without Adjustment
@@
Table B3: Peer Effects in Text Class Probabilities (Male) IV 2SLS without Adjustment
@@
Figure B1: Overview of Methods Pipeline and Results
Figure B2: Entry, Exit Dates, Length of Stay and Time to Recidivism
@@
Figure B3: Exchanges of Affirmations and Corrections (Count)
@@
Figure B4: Class Probabilities and Message Type in Male Units
@@
Figure B5: BERT Architecture
Notes: The architecture has been reproduced from the framework in Vaswani et al. [2017].
Figure B6: Predict Recidivism with and without text embeddings Aggregated by Receiver
@@
Figure B7: Predict Recidivism with and without class probabilities Aggregated by Receiver

### v2

@@ 6.2 Endogenous Network Formation
(c) Std. deviation of row means (Male)
(d) Std. deviation of row means (Female)
(e) Transitivity of clustering coef. (Male)
(f) Transitivity of clustering coef. (Female)
@@
We compare five estimators for four network models in terms of their ability to replicate observed salient features of the TC networks. These estimators/models are (1) tetra logit estimator (tet.logit) and the (2) joint fixed effect maximum likelihood estimator (jfe) in Graham, (2017) that allow for single latent variable (additive) and covariates, 3) Spectral embedding in RDPG model (Athreya et al.,, 2018) which allows for multivariate latent factors but no covariates (rdpg), 4) MLE in latent space model that allows for both multiplicative and additive latent factors but no observed covariates Hoff et al., (2002) (lsm) and finally 5) MLE in latent space model from Ma et al., (2020) that allows for both multiplicative and additive latent factors and observed covariates (lsm-cov).
Among the above-listed estimators that allow for observed covariates for network formation, we use the time difference (in days) between the residents’ entry dates as a node-pair level covariate. The entry dates of residents can be assumed to be random and not explained by their observed or unobserved characteristics, since they are determined by a complex function of court decisions and the availability of beds in TCs. However, the proximity of entry dates is an important determinant of whether residents will exchange messages since overlap in their time in the TC is a necessary condition to exchange messages. We do not use the proximity in exit dates or time in TC as a predictor of network formation, as the unobserved heterogeneity of residents likely impacts these variables, and they are not pre-determined to network formation.
Table 1: Cragg-Donald Weak Instruments Test
Gender Profile Type CD Statistic Critical Value
Female Sender 13.16 10.25
Receiver 13.61 10.25
Male Sender 14.68 10.25
Receiver 12.00 10.25
We compare the estimators in terms of their ability to replicate three widely observed structural network properties: modularity (Girvan and Newman,, 2002), standard deviation of row means, and transitivity or the clustering coefficient (Newman,, 2018). The steps deployed for comparison are as follows. Let us consider the network data for the male unit. We implement each estimator on the observed male unit network to estimate the respective fitted models. Next, we simulate 200 networks using these fitted models (200 for each estimator listed above). The structural properties over these simulated networks are then compared with the true network to assess how well the simulated data can replicate the truth. Figure 9 displays the performance. We observe that the latent space model that allows for both multiplicative and additive latent factors, along with covariates (proximity in entry dates), outperforms all other models for every property in the male unit. For the female unit, it performs the best for the standard deviation of row means and transitivity of the clustering coefficient, and at least as good as the other estimators for modularity.
@@ 6.3 Peer Effects in Text
Personal Growth Community Support Rule Violations
@@
This section provides our results on the peer effects model, where the outcome variables are the class probabilities generated by Zero-Shot classification. Our preferred specification involves estimating the network formation model using a latent space model, which allows for both multiplicative and additive latent variables and observed covariates (lsm-cov). This selection is based on the results obtained in the previous section, which show that the latent space model with covariates performs the best in replicating key structural properties of the observed male and female unit networks.
@@
Personal Growth Community Support Rule Violations
@@
As a first step, we assess the strength of the instruments for explaining the variability in the three endogenous regressors using the Cragg-Donald F statistic (CD statistic) and compare it with the Stock Yogo critical values (Stock and Yogo,, 2002). Table 1 shows the CD statistic is between 12 to 14.68 while the critical value at the 10% maximal IV relative bias is 10.25. This indicates the instruments are moderately strong. We note that maximal size-based Stock Yogo critical values are not available when the number of endogenous regressors is greater than 2 (three in our setup, one for each ALR), and hence we compare with only the relative bias critical values.
Tables 2 and 3 provide the results for IV 2SLS in text class probabilities after adjusting for the unobserved heterogeneity in the outcome model (Algorithm 1) for females and males, respectively. The dimension of the multiplicative factor in the latent space model is chosen using data-driven cross-validation methods for networks in Li et al., (2020). The choice of data-driven dimensions for the multiplicative factors is 14 for the female and 19 for the male units, respectively. In addition, the model for network formation contains one additive nodal latent variable and an observed edge-level covariate that captures the difference in the residents’ entry dates.
First, we discuss the diagonal elements in Table 2, which inform us about the spillovers of peer probabilities on female residents for the same class. There is evidence of a peer effect in all three classes relative to disruptive conduct at the 5% significance level. Moreover, there is also evidence of peer spillovers from personal growth and community support categories on rule violations category. We can interpret these results as follows, for example. An individual is more likely to send messages of peer community support if their peers are also sending messages of community support. As a comparison, Table A2.4 presents the results when unobserved heterogeneity in the outcome model and network formation are not accounted for.
@@
Overall, we find evidence of peer influence in the ALR of text class probabilities both when aggregated by the sender profiles and when aggregated by the receiver profiles. Interestingly, the magnitude and standard errors change considerably once we account for unobserved confounding in outcome and network formation using latent variables. This underscores the importance of controlling for latent confounding in peer effect studies. We also note from our results that the same category peer influence spillovers generally have a larger impact than across-category spillovers for both male and female units. Finally, we note that even though some of the values in our peer effect parameter matrix are greater than 1, the matrix $(I-D\otimes G)$ is invertible, which makes the structural multivariate peer effect model in Equation 1 stable (by Lemma 1 in Zhu et al., (2020)).
Robustness of Peer Effect Estimates:
Table 4: Network Permutation Test and Empirical P-Values
Personal Growth Community Support Rule Violations
Panel A: Female Unit (Sender Profiles)
Peer Personal Growth ALR 0.000 1.000 0.380
Peer Community Support ALR 0.880 0.040 0.500
Peer Rule Violations ALR 1.000 0.520 0.000
Panel B: Male Unit (Sender Profiles)
Peer Personal Growth ALR 0.600 0.240 0.860
Peer Community Support ALR 0.080 0.000 0.980
Peer Rule Violations ALR 0.100 0.580 0.000
Notes: Peers are randomly shuffled 50 times to estimate the empirical p-values for peer effects from network permutation. Diagonal elements (own peer effects) are the primary coefficients of interest.
We validate the peer effect estimates obtained in Section 6.3 by comparing them against the distribution of peer effects obtained by randomly permuting peers. Intuitively, the “significant” peer effects at conventional levels should be much higher in standardized absolute magnitude than this distribution since the peer effects with random peer assignment should be null. This is implemented by randomly shuffling the network connections, preserving each resident’s unique number of connections and the total number of messages exchanged. This process is repeated 50 times, yielding 50 simulated random networks. Once these networks are obtained, we repeat the steps outlined in Algorithm 1 to estimate the peer effects matrix. Note that, since the network changes with each simulation, we re-estimate the latent variables using our preferred latent-space model with covariates. The 50 networks and peer effect estimation can be used to construct the entire empirical distribution of the t-statistic of the peer effects parameter. The final step is to compute the empirical p-value as the proportion of test statistics that exceed the observed statistic value.
The empirical p-values for the observed test statistics of peer effects are reported in Table 4. The diagonal elements provide same-category peer effects, and the off-diagonal elements provide the cross-category effects. The table shows that each diagonal element has a very small empirical p-value (¡0.05). For the male unit, the observed t-statistics for peer effects on Community Support ($p=0.00$) and Rule Violations ($p=0.00$) are significant at conventional levels, whereas the peer effect on Personal Growth is not statistically significant ($p=0.60$). This is consistent with the observations from the main results in Table 3.
7 Discussion
The results in Section 6 show that there are substantial peer effects in language within the correctional units. The prediction analysis provides evidence that language profiles contain meaningful signals for predicting recidivism, especially in settings where conventional pre-entry covariates have limited predictive power. However, we consider a limitation of our analysis. For policy implications, one has to move beyond predictive power. To establish the value of language profiles for recidivism, it is necessary to estimate the causal link between language profiles and post-release outcomes.
The directed acyclic graph for this causal effect is provided in Figure 10. The primary objective is to estimate the direct effect of the $i^{th}$ resident’s language profile ($L_{i}$) on the $i^{\prime}s$ recidivism ($R_{i}$), conditional on the observed covariates $X_{i}$. However, the unobserved heterogeneity $U_{i}$ confounds this relationship. These unobservables can capture multiple factors such as behaviors, personality traits, motivation levels, and aspirations.
We estimate the linear relationship between $R_{i}$ and $L_{i}$ controlling for the observed covariates. Next, we do sensitivity analysis using two complementary methods Oster, (2019) and Cinelli and Hazlett, (2020) to get the bounds on how much omitted variable bias is needed relative to the observables to explain away the entire effect of $L_{i}$ on $R_{i}$.
Figure 10: Causal Impact of Own Language on Recidivism
$L_{i}$$R_{i}$$U_{i}$$X_{i}$$\beta$
First, we implement the approach outlined in Oster, (2019). Oster, (2019) defines the “coefficient of proportionality” $\delta$ that informs the extent of the selection on unobservable relative to the observed covariates needed to completely explain the estimated effect of $L_{i}$ on $R_{i}$. To calculate $\delta$, the researcher has to fix the value of $R_{\text{max}}$, which is the true $R^{2}$ if we were to regress $R_{i}$ on $L_{i},X_{i}$, and the unobservable $U_{i}$. For our analysis we set $R_{\text{max}}=1.3\tilde{R}$ using the recommendations in Oster, (2019) where $\tilde{R}$ is the $R^{2}$ from the regression of $R_{i}$ on the observables $L_{i},X_{i}$.
Table A2.6 provides the results of this analysis for the female sender, female receiver, male sender, and receiver in panels A–D, respectively. Column (1) reports the estimated $\beta$ from the regression of $R_{i}$ on $L_{i}$ controlling for the observed covariates (LSI-R, age, education, and race). In Column (2), we report the $\beta$ for the benchmark scenario where the strength of the selection on unobservables is equivalent to that of the selection on observed covariates. Column (3) reports the $\delta^{*}$ that will eliminate the effect of $L_{i}$ on $R_{i}$. Finally, the last column reports the $R^{2}$ from the model when $R_{i}$ is regressed on both the language profiles and the covariates.
For the female unit, all three language profile coefficients remain economically meaningful even when considering that the selection on unobservables is as strong as the selection on observables. For instance, in the female sender profiles, Community Support reduces recidivism by 2.4 percentage points (versus 3.7 in OLS), while Rule Violations increases it by 7.6 percentage points (versus 11.8 in OLS). We see similar patterns, but only for Community support and Rule Violation ALRs in the male unit. Lastly, the results in column (3) suggest that the selection on unobservables has to be in the range of 2.7–2.9 times the selection on observables to completely explain away the impact of language profiles on recidivism.
For the second bounding analysis presented in Table A2.7, we use the “sensmakr” method in Cinelli and Hazlett, (2020). Unlike our previous bounding analysis, sensemakr R package does not support simultaneous sensitivity analysis for multiple correlated treatments. Therefore, we construct a single language treatment variable, which we call the “positive language” variable. To do so, we add the positive zero-shot classes Personal Growth and Community Support to create a positive language class, and the negative zero-shot classes Rule Violations and Disruptive Conduct to create a negative class. We then take the log ratio of these two new classes to obtain our positive language variable.
From the Table A2.7, we see the Robustness Values $RV(q=1)$ ranges from 0.22 to 0.24. This indicates that, with the assumption of equal strength of association with the outcome and the treatment (in terms of partial $R^{2}$, i.e., explaining the residual variance), a confounder needs to explain 22-24% of the residual variation in both the outcome and the treatment in a model that also contains the observed covariates. A similar observation can be made from the column of $RV(\alpha=0.05)$ in terms of the strength of confounders needed to completely invalidate the statistical significance of the effects of positive language. Benchmarking against the observed covariate LSI-R, we see that the statistical significance of the effect of positive language will remain (for female senders) even if we have an unobserved confounder that is 3 times as strong as LSI-R in terms of association with both the outcome and the treatment.
The bounding analysis above provides an estimate of the strength of unobserved confounders required to invalidate our estimate. However, we note that to satisfactorily establish the causal impact of language profiles on recidivism, one needs either to implement a randomized controlled trial that creates random variation in language or to use natural experiments to tease out the causal effect. We see this as a future direction of research, building on our work on the underlying mechanisms of peer effects in language interactions in these facilities.
In conclusion, we develop a framework that combines AI-based measurements of unstructured text with econometric methods to investigate peer effects in language. This framework can be applied more generally, outside our application to recidivism, to study peer effects in contexts such as education and worker productivity, where written or spoken language exchange data are available. Our results also highlight the predictive ability of AI measurements to complement traditional covariates for predicting economic outcomes.
Appendix A1 Mathematical Appendix
@@ 7 Discussion
We can easily see that the conditional expectation of the second part, conditioning on $\bm{U}_{i}$ is (adapting the arguments from Johnsson and Moon, (2021))
@@
A1.2 Proof of Theorem 1
We derive the asymptotic distribution of the 2SLS estimator $\hat{\boldsymbol{\beta}}_{v,\text{IV}}$ in three steps similar to the structure in Johnsson and Moon, (2021).
@@
Now, let ${\cal F}_{Ni}:=\sigma\left(\bm{X}_{N},\bm{A},\bm{u}_{s},\bm{e}_{s};\,s\leq i\right)$ be a filtration. Then, $T_{Ni}:=\frac{1}{\sqrt{N}}\tilde{\bm{k}}_{i}^{\prime}\bm{Q}\tilde{\bm{e}}_{i}$ is a martingale difference adapted to ${\cal F}_{Ni}$. For simplicity, let $\bm{\iota}_{i}^{E}:=\bm{Q}\tilde{\bm{e}}_{i}$, which are i.i.d from $(0_{3p},\bm{Q}\Sigma_{\tilde{E}}\bm{Q}^{\prime})$. We check the sufficient conditions for Corollary 3.1 of Hall and Heyde, (2014) to apply the Martingale central limit theorem.
@@
    By Chen et al., (2023, equation 45 in Lemma A.8),
@@
A1.3 Proof of additional lemmas
A1.3.1 Proof of Lemma 2
This part of the proof follows similar logic as in Johnsson and Moon, (2021); however, since our terms are matrix-valued instead of vector-valued, our bounds involve the matrix spectral and Frobenius norms, as well as the matrix trace. We first prove the four results claimed in the lemma.
@@
    under Assumption 3 using the results of Lemmas 1 and 6. Next,
@@
A1.4 Additional Technical Lemmas
@@
Let Assumption 3 hold. Under the conditions on RDPG model in Rubin-Delanchy et al., (2022) and sparsity parameter $\rho_{N}=\omega(\frac{\log^{4}N}{N})$, we have
@@
and under the conditions on the sparse LSM model with sparsity parameter $\omega_{N}=\omega(N^{-1/2})$ in Li et al., (2023), we have
@@
By the results on Adjacency Spectral Embedding in RDPG model in Rubin-Delanchy et al., (2022), the last term is less than
@@
Temporarily write $\bm{W}^{(i)}:=\bm{\phi}^{L_{N}}(\bm{u}_{i})\bm{\phi}^{L_{N}}(\bm{u}_{i})^{\prime}$. By Lemma 5 of Johnsson and Moon, (2021), we have
@@
A1.5 Algorithm
Algorithm 1 Vectorized 2SLS with Network Endogeneity in MSAR
[…]
1:Step 1: Spectral Embedding (RDPG)
[…]
3:Step 2: Construct Peer Lag Matrices
[…]
6:Step 3: Latent Adjustment
[…]
10:Step 4: Vectorization
[…]
14:Step 5: IV Estimation and Covariance Calculation
[…]
Appendix A2 Additional Tables and Figures
Table A2.1: Summary Statistics
@@ B. Male Unit
Table A2.2: Monte Carlo Simulations RDPG Model
Y1 Equation Y2 Equation
N $D_{11}$ $D_{12}$ $D_{21}$ $D_{22}$
$h^{Y}(U)=U\Theta_{Y}$
100 1.349 0.284 1.662 0.383
300 1.057 0.224 1.041 0.217
500 0.471 0.098 0.472 0.100
$h^{Y}(U)=[U[,1],U[,2],U[,1]^{2},U[,2]^{2}]\Theta_{Y}$
100 0.638 0.218 0.777 0.279
300 0.256 0.090 0.157 0.055
500 0.205 0.073 0.113 0.039
$h^{Y}(U)=[U[,1],U[,2],U[,1]^{2},U[,2]^{2},U[,1]*U[,2]]\Theta_{Y}$
100 1.671 0.635 1.167 0.460
300 0.514 0.195 0.693 0.258
500 0.156 0.062 0.220 0.091
Table A2.3: Monte Carlo Simulations Latent Space Model with Covariates
Y1 Equation Y2 Equation
N $D_{11}$ $D_{12}$ $D_{21}$ $D_{22}$
$h^{Y}(U)=U\Theta_{Y}$
100 2.434 3.711 1.489 2.241
300 0.874 1.169 0.686 1.064
500 0.413 0.644 0.666 0.833
$h^{Y}(U)=[U[,1],U[,2],U[,1]^{2},U[,2]^{2}]\Theta_{Y}$
100 1.362 1.700 4.380 5.600
300 0.670 0.861 3.957 5.093
500 0.263 0.338 0.987 1.268
$h^{Y}(U)=[U[,1],U[,2],U[,1]^{2},U[,2]^{2},U[,1]*U[,2]]\Theta_{Y}$
100 1.820 2.261 3.256 4.055
300 1.120 1.423 1.933 2.478
500 0.605 0.769 0.838 1.072
Table A2.4: Peer Effects in Text Class Probabilities (Female) IV 2SLS without Adjustment
@@
Table A2.5: Peer Effects in Text Class Probabilities (Male) IV 2SLS without Adjustment
@@
Table A2.6: Sensitivity Analysis: Language Dimensions and Recidivism
Language Dimension $\beta$ (OLS) $\beta^{*}(\delta=1)$ $\delta^{*}$ $R^{2}_{full}$
Panel A: Female Unit, Sender Profiles
Personal Growth ALR -0.080 -0.051 2.782 0.093
Community Support ALR -0.037 -0.024 2.782 0.093
Rule Violations ALR 0.118 0.076 2.782 0.093
Panel B: Female Unit, Receiver Profiles
Personal Growth ALR -0.023 -0.015 2.888 0.115
Community Support ALR -0.106 -0.069 2.888 0.115
Rule Violations ALR 0.183 0.120 2.888 0.115
Panel C: Male Unit, Sender Profiles
Personal Growth ALR -0.006 -0.004 2.897 0.076
Community Support ALR -0.132 -0.087 2.897 0.076
Rule Violations ALR 0.048 0.031 2.897 0.076
Panel D: Male Unit, Receiver Profiles
Personal Growth ALR 0.010 0.006 2.891 0.075
Community Support ALR -0.137 -0.090 2.891 0.075
Rule Violations ALR 0.070 0.046 2.891 0.075
Notes: All models include three language dimensions jointly (Personal Growth, Community Support, Rule Violations) plus controls (LSI-R, age, education, race). $\beta$ (OLS) is the OLS coefficient with standard error in parentheses. $\beta^{*}(\delta=1)$ is the bounded estimate assuming selection on unobservables equals selection on observables. $\delta^{*}$ is the degree of selection on unobservables relative to selection on observables needed to completely eliminate the treatment effect of language on recidivism. $R^{2}_{full}$ is from the full model with all three language dimensions. $R^{2}_{restricted}$ (controls only) ranges from 0.015 to 0.017 across all models.
Table A2.7: Sensitivity to Unobserved Confounding: Positive Language
Unit Profile N $\beta$ Partial $R^{2}$ RV ($q=1$) RV ($\alpha=0.05$)
Panel A: Female Unit
Female Sender 982 -0.142 0.071 0.241 0.192
Female Receiver 982 -0.166 0.090 0.268 0.221
Panel B: Male Unit
Male Sender 1649 -0.144 0.064 0.229 0.191
Male Receiver 1649 -0.134 0.062 0.226 0.187
Notes: Positive language is log[(Personal Growth + Community Support) / (Rule Violations + Disruptive Conduct)]. All models include controls (LSI-R, age, education, race). Standard errors in parentheses. Partial $R^{2}$ is the partial $R^{2}$ of treatment with outcome. RV ($q=1$) is the percentage of residual variance an unobserved confounder must explain for both treatment and outcome to drive the effect to zero. RV ($\alpha=0.05$) is the robustness value to make the effect statistically insignificant at $\alpha=0.05$.
Figure A2.1: Entry, Exit Dates, Length of Stay and Time to Recidivism
@@
Figure A2.2: Exchanges of Affirmations and Corrections (Count)
@@
Figure A2.3: Class Probabilities and Message Type in Male Units
@@
Figure A2.4: BERT Architecture
Notes: The architecture has been reproduced from the framework in Vaswani et al., (2017).
Figure A2.5: Calibration Plots Predicted Probability and True Recidivism (Male)
a. Histogram True vs. Predicted Outcome
b. Calibration Plot
Notes: Five-fold cross-fitting is used to estimate the out-of-sample probabilities for residents in each fold.
Figure A2.6: Predict Recidivism with and without text embeddings Aggregated by Receiver
@@
Figure A2.7: Predict Recidivism with and without class probabilities Aggregated by Receiver
@@
Appendix A3 Zero-Shot Classification Label Selection
A principled, iterative, and reproducible LLM query approach is used to generate the input labels for zero-shot classification. We created 100 random samples of affirmations and corrections after appending the text messages of the male and female units. The random sample generation was stratified by message type (affirmations and corrections). In other words, in every sample about 500 messages were affirmations and 500 were corrections. Moreover, each sample was generated using a random seed to maintain independence across samples.
Each sample was submitted to the Claude API (Anthropic, claude-sonnet-4) with the following prompt:
“I want to do zero-shot text classification. I have a sample of 1000 texts from a therapeutic community in a correctional facility - short messages exchanged between residents (affirmations and corrections). I want to classify these into 2 positive and 2 negative behaviors. Based on these messages, suggest 4 class names (2 positive, 2 negative) that are mutually exclusive and relevant for predicting recidivism. Return ONLY the 4 class names, one per line, with positive classes first, then negative classes.
Messages: [1000 text messages]”
Across 100 iterations, the suggested labels showed strong convergence. Table A3.8 summarizes the clustering patterns observed.
Table A3.8: Summary of Suggested Class Labels Across 100 Iterations
Category Common Suggestions Selected Label
Positive 1 Program Engagement, Supportive Engagement, Community Support
Prosocial Leadership, Positive Leadership
Positive 2 Personal Growth, Personal Development, Personal Growth
Academic Achievement, Program Commitment
Negative 1 Rule Violations , Rule Violations
Rule Violation, Rule Breaking
Negative 2 Disrespectful Behavior, Disruptive Behavior, Disruptive Conduct
Disrespectful Conduct, Interpersonal Aggression
The label “Rule Violations” demonstrated the highest stability, appearing in approximately 95% of iterations. Positive behavior labels clustered into two distinct semantic groups: one centered on program engagement and prosocial behavior, and another focused on personal growth and achievement. Negative behavior labels similarly clustered around rule-breaking and interpersonal misconduct themes.
Figure A3.8: Zero Shot Classification
Based on this analysis, we selected four final labels for zero-shot classification: Personal Growth, Community Support, Rule Violations, and Disruptive Conduct. These labels are representative of the distribution of API-generated categories while maintaining mutual exclusivity and relevance for predicting recidivism.
Finally, to examine the semantic relationships among zero-shot classification labels, we embedded all candidate labels using BERT-base-uncased model and projected the 768-dimensional embeddings into a 2D space using UMAP (McInnes et al.,, 2018). UMAP projection used parameters n_neighbors=15n\_neighbors=15 and min_dist=0.1min\_dist=0.1 to preserve local neighborhood structure. Figure A3.8 shows the resulting semantic space of labels. Labels cluster into interpretable groups. Our pre-specified labels (shown in red) align well with API-generated negative-behavior labels, while positive labels form distinct clusters around concepts such as leadership, community support, program engagement, and personal growth.
Appendix A4 Multi-Head Self-Attention and transformer blocks
The query vector represents what the $k$-th token seeks, and the key vector indicates what each token offers others. The attention score between the $k$-th query and all $K$ keys is given by $\mathbf{s}_{k}^{(l,a)}=\left[s_{k1}^{(l,a)},s_{k2}^{(l,a)},\ldots,s_{kK}^{(l,a)}\right]^{\top}\in\mathbb{R}^{K}$. Each element in this vector is the scaled dot-product between the corresponding query and key vectors, $s_{kj}^{(l,a)}=\frac{\mathbf{q}_{k}^{(l,a)\top}\mathbf{k}_{j}^{(l,a)}}{\sqrt{d}}.$ The attention weights are then obtained by normalizing $s_{kj}^{(l,a)}$ using the softmax function as $m_{kj}^{(l,a)}=\frac{\exp(s_{kj}^{(l,a)})}{\sum_{q=1}^{K}\exp(s_{kq}^{(l,a)})}.$ Let $\mathbf{M}^{(l,a)}\in\mathbb{R}^{K\times K}$ be the attention weights matrix for all tokens in $\mathbf{z}$. The output matrix is given by $\mathbf{H}^{(l,a)}=\mathbf{M}^{(l,a)}\mathbf{V}^{(l,a)}\in\mathbb{R}^{K\times d}$. Next, the outputs from $A$ heads are combined as: $\mathbf{H}^{(l)}=\left[\mathbf{H}^{(l,1)},\mathbf{H}^{(l,2)},\ldots,\mathbf{H}^{(l,A)}\right]\in\mathbb{R}^{K\times H}$. The column-concatenated matrix $\mathbf{H}^{(l)}$ is then processed within the encoder block to generate the block’s final output $\mathbf{E}^{(l)}$ (Figure A2.4).